A helium-neon laser diffraction-limited light beam of wavelength 600 nm and diameter 3×10⁻³m is directed towards the surface of the moon 375×10³ km away.
To calculate the diameter of light patch produced on the moon.
1
Expert's answer
2019-12-13T10:11:50-0500
"\Delta \theta=\frac{2.44\lambda}{D}=\frac{2.44(600\cdot 10^{-9})}{3\cdot 10^{-3}}=0.000488"

"W=L\Delta \theta=(375000000)(0.000488)=183000\ m=183\ km"